# A Survey of 3D Reconstruction with Event Cameras

### v2

@@ IV-B Learning-Based Methods with Traditional Output
Without pipeline and priors: However, recent methods aim to eliminate the pipeline and priors. In 2023, Chen et al. [95] proposed E2V, which employs a modified ResNet-152 and a U-Net 3D decoder to directly predict dense 3D voxel grids from monocular event frames, achieving event-based 3D reconstruction without external priors. In 2025, Xu et al. [27] extended E2V by introducing a novel event representation, Sobel Event Frame, and an optimal binarisation strategy for event-based 3D reconstruction. By enhancing E2V with Efficient Channel Attention [118, 119], their method significantly improved reconstruction quality.
IV-C Methods with Neural Radiance Fields
@@
Figure 10: Overview of methods with monocular event cameras that produce Neural Radiance Fields. (a) The method structure mainly modelling brightness difference, $\Delta\hat{L}_{k}$, from Klenk et al. [35]. (b) The method structure that also models brightness change rate, $\frac{\partial}{\partial t}\log\hat{L}(\mathbf{u},t_{\text{sam}})$, from Low et al. [120].
TABLE V: Monocular event camera: NeRF-based Methods
@@
Klenk et al. [35] E-NeRF 2023-01 Event stream EAF ESIM [105]
Hwang et al. [121] Ev-NeRF 2023-03 Event stream EAF IJRR [122], HQF [123]
Rudnev et al. [96] EventNeRF 2023-06 RGB bayer event stream EAF Self-collected [96]
Low et al. [120] Robust e-NeRF 2023-10 Event stream Event-by-event TUM-VIE [124]
Bhattacharya et al. [125] EvDNeRF 2024-01 Event stream EAF Self-collected [125]
Wang et al. [109] NeRF(Enhanced) 2024-05 Event stream EAF PAEv3D [109]
Low et al. [126] Deblur e-NeRF 2024-09 Event stream Event-by-event EDS [127]
Feng et al. [128] AE-NeRF 2025-01 Event stream Event-by-event TUM-VIE [124]
Wang et al. [129] SaENeRF 2025-04 Event stream EAF Rudnev et al. [96]
@@
Figure 11: Visual comparison of several NeRF-based methods using monocular event cameras, adapted from Feng et al. [128], under CC BY-NC-SA 4.0. (a) Result from the monocular deep learning method E2VID combined with NeRF [104]. (b) Ev-NeRF method [121] (c) Robust e-NeRF method [120]. (d) AE-NeRF method [128].
@@
Compared to geometry-based or deep learning-based methods, NeRF-based methods avoid explicit feature extraction or geometry construction pipelines. Instead, they rely on end-to-end optimisation guided by photometric or event-based supervision to recover implicit scene structure. NeRF-based methods typically use the Event Accumulate Frame as the main input representation of event streams and are generally designed for grayscale 3D reconstruction. Particularly, some methods achieve colour 3D reconstruction using only the event stream as input. Low et al. [120, 126] restore colour using gamma correction. Feng et al. [128] adopt a learning-based colour correction method. Wang et al. [129] leverage NeRF’s volume rendering and colour modelling to perform self-supervised colour 3D reconstruction without RGB image supervision. Table V and Figure 10 provide an overview of these methods, and Figure 11 provides a visualisation example of some methods. Based on the physical modelling approach, we categorise them into the following two types:
Brightness difference modelling: Some relatively early methods supervise NeRF training only by matching the brightness difference between discrete time points, typically over fixed intervals or adjacent event pairs. These methods do not explicitly model the brightness change rate. In 2023, Hwang et al. [121] proposed Ev-NeRF, which aggregates events through multi-view consistency. By estimating the volumetric density and radiance, the method achieves high-quality depth reconstruction and novel view synthesis. In the same year, Rudnev et al. [96] proposed EventNeRF, which employs event-based volumetric rendering in a self-supervised manner to reconstruct high-quality 3D structures and synthesise new views. Later that year, Klenk et al. [35] proposed E-NeRF, which utilises an event-triggered brightness model along with a no-event loss to enable dense reconstruction. In 2024, Bhattacharya et al. [125] proposed EvDNeRF, the first dynamic event-based NeRF. EvDNeRF uses an event-triggered brightness model and a varied batching strategy to achieve high-fidelity dynamic reconstruction. In 2025, Wang et al. [129] proposed SaENeRF, which normalizes radiance variations based on accumulated event polarities and introduces zero-event regularization losses, enabling artifact-suppressed and photorealistic 3D reconstruction directly from event streams.
@@
Figure 12: Overview of methods with monocular event cameras that produce 3D Gaussian Splatting. (a) Processing Structure from [130], including Temporal-to-Intensity Mapping to convert exposure events into intensity images, which yield camera trajectories and a sparse point cloud for 3DGS training. (b) Processing Structure from [98], including physical optimisation by aligning the rendered brightness changes with event-based measurements over adaptive temporal windows. (c) End-to-end learning structure from [131].
TABLE VI: Monocular event camera: 3DGS-based Methods
@@
Wang et al. [131] EvGGS 2024-05 Event stream Voxel Grid, EAF Ev3DS [131]
Zhang et al. [132] Elite-EvGS 2024-09 Event stream EAF Rudnev et al. [96]
Wu et al. [133] Ev-GS 2024-09 Event stream EGA Self-collected [133]
Huang et al. [134] IncEventGS 2024-10 Event stream EAF TUM-VIE [124]
Yin et al. [130] E-3DGS 2024-10 Event stream Event-by-event EME-3D [130]
Yura et al. [97] EventSplat 2024-12 RGB bayer event stream EAF EDS [127], TUM-VIE [124]
Jeong et al. [135] EOGS 2024-12 Event stream EAF EDS [127]
Zahid et al. [98] E-3DGS 2025-02 RGB bayer event stream EAF E-3DGS dataset [98]
Brightness change rate modelling: Some later methods incorporate loss functions that explicitly model the brightness change rate, defined as the brightness difference divided by the corresponding time interval. This enables finer temporal resolution, improves robustness to non-uniform motion, and better aligns with the physical characteristics of event triggering. In 2023, Low et al. [120] proposed Robust e-NeRF. This method introduces a more realistic event generation model and two normalised loss functions: one based on contrast-normalised difference and another on target-normalised temporal gradient. As a result, the model no longer requires known contrast thresholds or explicit event accumulation strategies, enabling robust self-supervised learning. In 2024, Wang et al. [109] proposed Physical Priors Augmented EventNeRF, which incorporates motion and geometric priors and adopts a density-guided patch sampling strategy to enhance structural representation. Also in 2024, Low et al. [126] proposed Deblur e-NeRF, which models pixel bandwidth to account for event motion blur and introduces a threshold-normalised total variation loss, enabling robust 3D reconstruction directly from motion-blurred event streams. In 2025, Feng et al. [128] proposed AE-NeRF, which integrates pose correction and hierarchical architecture to reconstruct NeRFs from sparse and asynchronous event streams accurately. To enhance visual quality, they use a colour correction network to recover RGB images from log-radiance.
IV-D Methods with 3D Gaussian Splatting
@@
Gaussian Splatting-based methods represent scenes using a set of 3D Gaussian primitives with learnable properties. Each primitive has learnable attributes, including position, covariance, and colour. Similar to NeRF-based methods, most Gaussian Splatting-based methods also use the Event Accumulation Frame as the main input representation of event streams. Table VI and Figure 12 provide an overview of these methods. Based on the source of supervision and structural guidance, we categorise these methods into the following three types:
@@
Reconstructing images as guidance: Some methods use image frames reconstructed from event streams as a source of visual guidance or as an initial structural prior to facilitate the learning process. In 2024, Zhang et al. [132] proposed Elite-EvGS, which distils prior knowledge from off-the-shelf event-to-video (E2V) models [136, 137]. It uses E2V-generated frames to initialise a coarse 3DGS model and then progressively incorporates raw events to refine scene details through event supervision. In the same year, Yin et al. [130] proposed E-3DGS, which uses a novel temporal-to-intensity mapping as visual guidance to facilitate 3D representation learning. The method also incorporates an event-type-specific supervision strategy and a hybrid optimisation approach. Also in 2024, Yura et al. [97] proposed EventSplat, which combines E2V guided SfM [113] initialisation and spline interpolation. It recovers continuous camera trajectories and achieves high-quality 3D reconstruction.
@@
Geometric optimisation: Some methods focus on explicit pose refinement and joint recovery of geometric structure. In 2024, Huang et al. [134] proposed IncEventGS, which follows a SLAM-inspired framework [138] and jointly estimates camera trajectory and 3D structure from event streams. In 2025, Zahid et al. [98] proposed E-3DGS, which uses frustum-based initialisation to generate an initial Gaussian point cloud. It extracts multi-scale structure and detail using adaptive event windows, and refines camera poses through an event loss to improve trajectory accuracy.
@@
End-to-end learning: Some methods focus on directly learning the final outputs from end-to-end learning structure. In 2024, Wang et al. [131] proposed EvGGS that connects depth estimation, intensity reconstruction, and 3D Gaussian parameter regression in a collaborative learning framework. The joint training improves reconstruction accuracy and rendering efficiency. In the same year, Wu et al. [133] proposed Ev-GS that combines neuromorphic imaging with 3DGS, modelling logarithmic brightness changes and enabling fast convergence using only event-based supervision.
V Multimodal Methods with Event Cameras
@@
Multimodal 3D reconstruction refers to methods that combine event camera data with other sensing modalities [139, 140, 141]. These methods aim to enhance reconstruction performance by leveraging the complementary strengths of each modality: the high temporal resolution and low latency of event cameras, and the rich spatial or appearance information provided by other sensors. Compared to stereo and monocular setups, multimodal systems are more flexible in input configurations. They often achieve higher reconstruction accuracy and robustness in low-light, high-speed, or motion-blurred environments.
@@
Early research in this direction typically combines event cameras with active sensing devices, such as structured light projectors [34], which enable high-speed depth recovery through time-coded patterns, and RGB-depth (RGB-D) cameras [140], which provide readily available depth data for event fusion. Many recent works fuse asynchronous events with frame-based RGB to construct coloured 3D models, or integrate them into advanced neural rendering frameworks such as Neural Radiance Fields and 3D Gaussian Splatting [139, 142]. These approaches enable photorealistic, temporally-aware 3D reconstruction under challenging real-world conditions.
@@
Based on the output representation and modelling strategy, multimodal methods can be broadly classified into three categories: (V-A) methods with traditional 3D outputs such as point clouds, (V-B) methods with Neural Radiance Fields, and (V-C) methods with 3D Gaussian Splatting.
V-A Multimodal Methods with Traditional Outputs
Multimodal methods with traditional outputs recover geometric structures by combining event data with complementary sensors such as structured light or RGB-D cameras. These systems fuse the temporal precision of events with the spatial density of external inputs, enabling robust reconstruction under challenging conditions. Table VII and Figure 13 provide an overview of these methods. As most of them rely on self-collected datasets, datasets are not included in Table VII. Based on the type of complementary modality integrated with event data, we categorise these methods into the following two types:
Combine with Structured light: Structured light is an active 3D sensing technique that projects coded patterns onto a surface and reconstructs depth via triangulation [143]. When integrated with event cameras, structured light systems can achieve robust, high-speed depth sensing under challenging conditions. Recent works have introduced several types of encoding and fusion strategies to use both spatial and temporal features of event streams.
In 2015, Matsuda et al. [141] proposed MC3D, one of the earliest event-based structured light systems. The method correlates laser scan timing with event timestamps, achieving high-precision depth reconstruction with per-pixel single-shot efficiency. In 2018, Leroux et al. [34] proposed event-based structured light systems by projecting frequency-tagged light patterns onto the scene. Each spatial region is illuminated with a unique modulation frequency, and the event camera decodes depth by associating detected frequencies with pixel locations. In 2021, Huang et al. [144] combined structured light projection with digital image correlation (DIC) [145, 146] for high-speed scanning. Also in 2021, Muglikar et al. [147] proposed ESL, which estimates depth by maximising spatio-temporal consistency between a laser projector and an event camera. By processing events in local space-time regions, their method improves robustness to noise. ESL accurately estimates depth in challenging scenes but is not real-time due to high computational cost. In 2023, Xiao et al. [148] employed alternating binary speckle patterns and DIC-based stereo matching for fast and accurate reconstruction. Also in 2023, Fu et al. [149] introduced spatio-temporal coding (STC) with an enhanced matching scheme for improved stereo robustness. In 2024, Li et al. [150] proposed eFPSL, using time-frequency analysis to extract high-SNR fringe maps from events and an event-count-based shadow mask to reduce errors.
@@ V-A Multimodal Methods with Traditional Outputs
Figure 13: Overview of multimodal methods with traditional outputs. (a) Schematic of event cameras working with structured light projector from Xiao et al. [148]. (b) Schematic of a stereo setup of an RGB-D camera and a monocular event camera from Zuo et al. [151].
@@
TABLE VII: Event cameras in multimodal system: Methods with traditional outputs
@@
Weikersdorfer et al. [140] 2014 RGB-D, Event stream Trajectory Event-by-event Point cloud (✗)
Matsuda et al. [141] 2015 Structured light, Event stream - Event-by-event 3D depth map (✓)
@@
Leroux et al. [34] 2018 Structured light, Event stream Pose Time surface Point cloud (✓)
@@
Huang et al. [144] 2021 Structured light, Event stream Pose Event-by-event Point cloud (✓)
Muglikar et al. [147] 2021 Structured light, Event stream - Event-by-event 3D depth map (✓)
Zuo et al. [151] 2022 RGB-D, Event stream Trajectory Time surface Point cloud (✗)
@@
Xiao et al. [148] 2023 Structured light, Event stream Pose Event frame Point cloud (✓)
@@
Fu et al. [149] 2023 Structured light, Event stream Pose Time surface 3D depth map (✗)
Li et al. [150] 2024 Structured light, Event stream Pose Event-by-event 3D depth map (✗)
Combine with RGB-D: Some methods fuse event data with RGB-D sensors for improved scene understanding from depth and RGB information. In 2014, Weikersdorfer et al. [140] proposed EB-SLAM-3D, a novel event-based 3D SLAM algorithm using a D-eDVS, enabling low-power, low-latency mapping with a sparse voxel grid at 20× real-time speed. In 2022, Zuo et al. [151] proposed DEVO, combining time surface maps from events and depth supervision from a calibrated sensor. Their system performs semi-dense 3D-2D edge alignment to estimate poses and incrementally build point clouds under fast motion and poor lighting.
V-B Multimodal Methods with Neural Radiance Fields
@@
Figure 14: Overview of multimodal methods with Neural Radiance Fields. (a) The method structure from Qi et al. that performs static scene reconstruction [152]. (b) The method structure from Rudnev et al. that performs dynamic scene reconstruction [153].
Multimodal Event-based NeRF methods combine event streams with RGB images to achieve high-quality coloured 3D reconstruction, leveraging the temporal precision of events and the rich appearance information of RGB frames to enhance structural accuracy and visual fidelity under motion blur and low-light conditions. Table VIII and Figure 14 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
Static scene reconstruction: Some methods focus on static scene reconstruction by integrating event supervision with blurry RGB images [152, 154, 155, 156], or by designing physically inspired mechanisms [157, 158] to improve NeRF training stability and output sharpness. In 2023, Qi et al. [152] proposed E²NeRF, which introduces a blur rendering loss and an event rendering loss to improve NeRF training under motion blur and low-light settings. In 2024, Cannici et al. [154] proposed Ev-DeblurNeRF, which uses the Event Double Integral (EDI) [159] and a learnable event camera response function (eCRF) to reconstruct sharp NeRFs. It performs well under severe motion blur. Later that year, Qi et al. [155] proposed E3NeRF, which combines blurry images and events with spatial-temporal attention [119]. It introduces an event-enhanced rendering loss to guide learning and improves robustness in non-uniform motion and low-light scenes. In the same year, Tang et al. [156] proposed LSE-NeRF, which models sensor response differences with per-time embeddings and event-based reflectance mapping to recover high-quality NeRFs without strict alignment. Later, Li et al. [160] proposed BeNeRF, which jointly reconstructs photorealistic 3D scenes and estimates camera motion from a single blurry RGB image and event stream, using an event rendering loss and B-spline trajectory representation for end-to-end optimisation. Around the same time, Qi et al. [157] proposed EBAD-NeRF, which introduces event-driven bundle adjustment. It jointly optimises NeRF and camera poses using an intensity-change event loss and a photometric blur loss. Also in 2024, Chen et al. [158] proposed Event-ID, which is the first event-driven framework for intrinsic decomposition. It combines an event-based reflectance model and a multi-view strategy to recover geometry, materials, and lighting under extreme conditions. In 2025, Chen et al. [161] proposed EvHDR-NeRF, which models a radiance-based relationship that accounts for exposure time and the camera response function (CRF), enabling HDR 3D reconstruction from single-exposure LDR images and event streams.
@@
TABLE VIII: Event cameras in multimodal system: NeRF-based methods

### v3

@@ 4.2.2 Without pipeline and priors
However, recent methods aim to eliminate the pipeline and priors. In 2023, Chen et al. [[Dense]] proposed E2V, which employs a modified ResNet-152 and a U-Net 3D decoder to directly predict dense 3D voxel grids from monocular event frames, achieving event-based 3D reconstruction without external priors. In 2025, Xu et al. [[Towards]] extended E2V by introducing a novel event representation, Sobel Event Frame, and an optimal binarisation strategy for event-based 3D reconstruction. By enhancing E2V with Efficient Channel Attention [[Towards-21], guo2022attention], their method significantly improved reconstruction quality.
4.3 Methods with Neural Radiance Fields
@@
Figure 10: Overview of methods with monocular event cameras that produce Neural Radiance Fields. (a) The method structure mainly modelling brightness difference, $\Delta\hat{L}_{k}$, from Klenk et al. (b) The method structure that also models brightness change rate, $\frac{\partial}{\partial t}\log\hat{L}(\mathbf{u},t_{\text{sam}})$, from Low et al.
Table 5: Monocular event camera: NeRF-based Methods
@@
Klenk et al. [[N3]] E-NeRF 2023-01 Event stream EAF ESIM [[esim]]
Hwang et al. [[N1]] Ev-NeRF 2023-03 Event stream EAF IJRR [IJRR], HQF [HQF]
Rudnev et al. [[N2]] EventNeRF 2023-06 RGB bayer event stream EAF Self-collected [[N2]]
Low et al. [[N13]] Robust e-NeRF 2023-10 Event stream Event-by-event TUM-VIE [TUM-VIE]
Bhattacharya et al. [[N4]] EvDNeRF 2024-01 Event stream EAF Self-collected [[N4]]
Wang et al. [[N7]] NeRF(Enhanced) 2024-05 Event stream EAF PAEv3D [[N7]]
Low et al. [[XN2]] Deblur e-NeRF 2024-09 Event stream Event-by-event EDS [EDS]
Feng et al. [[N8]] AE-NeRF 2025-04 Event stream Event-by-event TUM-VIE [TUM-VIE]
Wang et al. [[XN4]] SaENeRF 2025-04 Event stream EAF Rudnev et al. [[N2]]
@@
Figure 11: Visual comparison of several NeRF-based methods using monocular event cameras, adapted from Feng et al., under $C$ BY-NC-SA 4.0. (a) Result from the monocular deep learning method E2VID combined with NeRF. (b) Ev-NeRF method. (c) Robust e-NeRF method. (d) AE-NeRF method.
@@
Compared to geometry-based or deep learning-based methods, NeRF-based methods avoid explicit feature extraction or geometry construction pipelines. Instead, they rely on end-to-end optimisation guided by photometric or event-based supervision to recover implicit scene structure. NeRF-based methods typically use the Event Accumulate Frame as the main input representation of event streams and are generally designed for grayscale 3D reconstruction. Particularly, some methods achieve colour 3D reconstruction using only the event stream as input. Low et al. [[N13], [XN2]] restore colour using gamma correction. Feng et al. [[N8]] adopt a learning-based colour correction method. Wang et al. [[XN4]] leverage NeRF’s volume rendering and colour modelling to perform self-supervised colour 3D reconstruction without RGB image supervision. Table 5 and Figure 10 provide an overview of these methods, and Figure 11 provides a visualisation example of some methods. Based on the physical modelling approach, we categorise them into the following two types:
@@
Figure 12: Overview of methods with monocular event cameras that produce 3D Gaussian Splatting. (a) Processing Structure from Yin et al., including Temporal-to-Intensity Mapping to convert exposure events into intensity images, which yield camera trajectories and a sparse point cloud for 3DGS training. (b) Processing Structure from Zahid et al., including physical optimisation by aligning the rendered brightness changes with event-based measurements over adaptive temporal windows. (c) End-to-end learning structure from Wang et al.
4.3.1 Brightness difference modelling
Some relatively early methods supervise NeRF training only by matching the brightness difference between discrete time points, typically over fixed intervals or adjacent event pairs. These methods do not explicitly model the brightness change rate. In 2023, Hwang et al. [[N1]] proposed Ev-NeRF, which aggregates events through multi-view consistency. By estimating the volumetric density and radiance, the method achieves high-quality depth reconstruction and novel view synthesis. In the same year, Rudnev et al. [[N2]] proposed EventNeRF, which employs event-based volumetric rendering in a self-supervised manner to reconstruct high-quality 3D structures and synthesise new views. Later that year, Klenk et al. [[N3]] proposed E-NeRF, which utilises an event-triggered brightness model along with a no-event loss to enable dense reconstruction. In 2024, Bhattacharya et al. [[N4]] proposed EvDNeRF, the first dynamic event-based NeRF. EvDNeRF uses an event-triggered brightness model and a varied batching strategy to achieve high-fidelity dynamic reconstruction. In 2025, Wang et al. [[XN4]] proposed SaENeRF, which normalizes radiance variations based on accumulated event polarities and introduces zero-event regularization losses, enabling artifact-suppressed and photorealistic 3D reconstruction directly from event streams.
4.3.2 Brightness change rate modelling
Some later methods incorporate loss functions that explicitly model the brightness change rate, defined as the brightness difference divided by the corresponding time interval. This enables finer temporal resolution, improves robustness to non-uniform motion, and better aligns with the physical characteristics of event triggering. In 2023, Low et al. [[N13]] proposed Robust e-NeRF. This method introduces a more realistic event generation model and two normalised loss functions: one based on contrast-normalised difference and another on target-normalised temporal gradient. As a result, the model no longer requires known contrast thresholds or explicit event accumulation strategies, enabling robust self-supervised learning. In 2024, Wang et al. [[N7]] proposed Physical Priors Augmented EventNeRF, which incorporates motion and geometric priors and adopts a density-guided patch sampling strategy to enhance structural representation. Also in 2024, Low et al. [[XN2]] proposed Deblur e-NeRF, which models pixel bandwidth to account for event motion blur and introduces a threshold-normalised total variation loss, enabling robust 3D reconstruction directly from motion-blurred event streams. In 2025, Feng et al. [[N8]] proposed AE-NeRF, which integrates pose correction and hierarchical architecture to reconstruct NeRFs from sparse and asynchronous event streams accurately. To enhance visual quality, they use a colour correction network to recover RGB images from log-radiance.
4.4 Methods with 3D Gaussian Splatting
Table 6: Monocular event camera: 3DGS-based Methods
@@ 4.4 Methods with 3D Gaussian Splatting
Wang et al. [[G9]] EvGGS 2024-07 Event stream Voxel Grid, EAF Ev3DS [[G9]]
Wu et al. [[G6]] Ev-GS 2024-09 Event stream EGA Self-collected [[G6]]
Jeong et al. [[XG2]] EOGS 2024-12 Event stream EAF EDS [EDS]
Han et al. [[G7]] Event-3DGS 2024-12 Event stream Event-by-event DeepVoxels [[DeepVoxels]]
Zahid et al. [[G11]] E-3DGS 2025-03 RGB bayer event stream EAF E-3DGS dataset [[G11]]
Yin et al. [[G12]] E-3DGS 2025-05 Event stream Event-by-event EME-3D [[G12]]
Zhang et al. [[G10]] Elite-EvGS 2025-05 Event stream EAF Rudnev et al. [[N2]]
Yura et al. [[G5]] EventSplat 2025-06 RGB bayer event stream EAF EDS [EDS], TUM-VIE [TUM-VIE]
Huang et al. [[G8]] IncEventGS 2025-06 Event stream EAF TUM-VIE [TUM-VIE]
@@
Gaussian Splatting-based methods represent scenes using a set of 3D Gaussian primitives with learnable properties. Each primitive has learnable attributes, including position, covariance, and colour. Similar to NeRF-based methods, most Gaussian Splatting-based methods also use the Event Accumulation Frame as the main input representation of event streams. Table 6 and Figure 12 provide an overview of these methods. Based on the source of supervision and structural guidance, we categorise these methods into the following three types:
4.4.1 Reconstructing images as guidance
@@
Some methods use image frames reconstructed from event streams as a source of visual guidance or as an initial structural prior to facilitate the learning process. In 2025, Yin et al. [[G12]] proposed E-3DGS, which uses a novel temporal-to-intensity mapping as visual guidance to facilitate 3D representation learning. The method also incorporates an event-type-specific supervision strategy and a hybrid optimisation approach. Later, Zhang et al. [[G10]] proposed Elite-EvGS, which distils prior knowledge from off-the-shelf event-to-video (E2V) models [[G10-21], [G10-4]]. It uses E2V-generated frames to initialise a coarse 3DGS model and then progressively incorporates raw events to refine scene details through event supervision. Later that year, Yura et al. [[G5]] proposed EventSplat, which combines E2V guided SfM [[MD2-17]] initialisation and spline interpolation. It recovers continuous camera trajectories and achieves high-quality 3D reconstruction.
4.4.2 Geometric optimisation
@@
Some methods focus on explicit pose refinement and joint recovery of geometric structure. In 2025, Zahid et al. [[G11]] proposed E-3DGS, which uses frustum-based initialisation to generate an initial Gaussian point cloud. It extracts multi-scale structure and detail using adaptive event windows, and refines camera poses through an event loss to improve trajectory accuracy. In the same year, Huang et al. [[G8]] proposed IncEventGS, which follows a SLAM-inspired framework [[G8-20]] and jointly estimates camera trajectory and 3D structure from event streams.
@@
End-to-end learning: Some methods focus on directly learning the final outputs from end-to-end learning structure. In 2024, Wang et al. [[G9]] proposed EvGGS that connects depth estimation, intensity reconstruction, and 3D Gaussian parameter regression in a collaborative learning framework. The joint training improves reconstruction accuracy and rendering efficiency. In the same year, Wu et al. [[G6]] proposed Ev-GS that combines neuromorphic imaging with 3DGS, modelling logarithmic brightness changes and enabling fast convergence using only event-based supervision. Later, Han et al. [[G7]] proposed Event-3DGS, which introduces a high-pass filter-based photovoltage estimation module to reduce noise and enhance robustness effectively.
5 Multimodal Methods with Event Cameras
@@
Multimodal 3D reconstruction refers to methods that combine event camera data with other sensing modalities [[N9], [XDMT3], [XDMT1]]. These methods aim to enhance reconstruction performance by leveraging the complementary strengths of each modality: the high temporal resolution and low latency of event cameras, and the rich spatial or appearance information provided by other sensors. Compared to stereo and monocular setups, multimodal systems are more flexible in input configurations. They often achieve higher reconstruction accuracy and robustness in low-light, high-speed, or motion-blurred environments.
@@
Early research in this direction typically combines event cameras with active sensing devices, such as structured light projectors [[DMT1]], which enable high-speed depth recovery through time-coded patterns, and RGB-depth (RGB-D) cameras [[XDMT3]], which provide readily available depth data for event fusion. Many recent works fuse asynchronous events with frame-based RGB to construct coloured 3D models, or integrate them into advanced neural rendering frameworks such as Neural Radiance Fields and 3D Gaussian Splatting [[N9], [G2]]. These approaches enable photorealistic, temporally-aware 3D reconstruction under challenging real-world conditions.
@@
Based on the output representation and modelling strategy, multimodal methods can be broadly classified into three categories: (5.1) methods with traditional 3D outputs such as point clouds, (5.2) methods with Neural Radiance Fields, and (5.3) methods with 3D Gaussian Splatting.
5.1 Multimodal Methods with Traditional Outputs
@@
Figure 13: Overview of multimodal methods with traditional outputs. (a) Schematic of event cameras working with structured light projector from Xiao et al. (b) Schematic of a stereo setup of an RGB-D camera and a monocular event camera from Zuo et al.
@@
Table 7: Event cameras in multimodal system: Methods with traditional outputs
@@
Weikersdorfer et al. [[XDMT3]] 2014 RGB-D, Event stream Trajectory Event-by-event Point cloud (✗)
Matsuda et al. [[XDMT1]] 2015 Structured light, Event stream - Event-by-event 3D depth map (✓)
@@
Leroux et al. [[DMT1]] 2018 Structured light, Event stream Pose Time surface Point cloud (✓)
@@
Huang et al. [[DMT2]] 2021 Structured light, Event stream Pose Event-by-event Point cloud (✓)
Muglikar et al. [[XDMT2]] 2021 Structured light, Event stream - Event-by-event 3D depth map (✓)
Zuo et al. [[DMT3]] 2022 RGB-D, Event stream Trajectory Time surface Point cloud (✗)
@@
Xiao et al. [[DMT4]] 2023 Structured light, Event stream Pose Event frame Point cloud (✓)
@@
Fu et al. [[DMT5]] 2023 Structured light, Event stream Pose Time surface 3D depth map (✗)
Li et al. [[DMT6]] 2024 Structured light, Event stream Pose Event-by-event 3D depth map (✗)
Multimodal methods with traditional outputs recover geometric structures by combining event data with complementary sensors such as structured light or RGB-D cameras. These systems fuse the temporal precision of events with the spatial density of external inputs, enabling robust reconstruction under challenging conditions. Table 7 and Figure 13 provide an overview of these methods. As most of them rely on self-collected datasets, datasets are not included in Table 7. Based on the type of complementary modality integrated with event data, we categorise these methods into the following two types:
@@
Figure 14: Overview of multimodal methods with Neural Radiance Fields. (a) The method structure from Qi et al. that performs static scene reconstruction. (b) The method structure from Rudnev et al. that performs dynamic scene reconstruction.
@@
Table 8: Event cameras in multimodal system: NeRF-based methods